# Evidence for itinerant electron-local moment interaction in Li-doped -MnTe Tingjun Zhang

### v1

@@ II.3 Inelastic neutron scattering
INS measurements on Li-doped -MnTe were carried out using the TOF spectrometer MAPS at the ISIS Neutron and Muon Source, Rutherford Appleton Laboratory [ewings2019MAPS]. The sample, with a mass of approximately 2.2 g, was obtained from the same large single crystal used for the TOPAZ experiment and was mounted in the [, 0, ] scattering plane. Data were collected at K with an incident energy of meV under continuous rotation about the vertical axis of the sample. During the measurements, the accelerator operated with a synchrotron current of 152 A on target station 1. Measurements were carried out in two modes: high-resolution (HR) mode with a Fermi chopper frequency of Hz and a sample rotation range of 60 degrees, and high-flux (HF) mode with Hz and a rotation range of 120 degrees. Each mode was allocated approximately 60 hours of beamtime, evenly distributed over the full angle range. The data were reduced into .nxspe files using a step of 0.5 degrees. Data from the HF mode are presented in Figs. 2(a)(c)–(e), 3(b)(d)(f)–(i), 4, and S2, while data from the HR mode are shown in Figs. 2(c)–(e), S3, and S4.
Data reduction was performed using MANTID [arnold2014mantid], Horace [ewings2016horace], and SHIVER [shiver].
@@
II.4 Absolute unit normalization
INS data were normalized to absolute units during the data reduction process by measuring a standard vanadium sample of known mass [10.1063/1.4818323]. The vanadium measurements were performed in the same sample environment and with the same chopper and instrument settings as the sample measurements.
II.5 Linear spin wave theory (LSWT) Calculations
The spin Hamiltonian was determined from the INS spectra using linear spin wave theory (LSWT) implemented in the SpinW package [toth2015spinw]. For both compounds, a Heisenberg spin Hamiltonian with a single-ion anisotropy term was employed, expressed as
where indexes the Mn2+ ions, is the spin operator at site , and denotes its component. represents the exchange interaction between spins at sites and , with the summation taken over spin pairs. In this coordinate system, the -axis corresponds to the crystallographic -axis. The anisotropy term is positive for easy-plane anisotropy, as in the case of pure -MnTe, and negative for Ising-type anisotropy, as observed in Li-doped -MnTe. To obtain the spin wave dispersion relation , we performed Gaussian fitting to extract the energy center and its half-width at half-maximum (HWHM) at each wave vector along high-symmetry directions.
@@ II.5 Linear spin wave theory (LSWT) Calculations
II.6 Electrical transport measurements
@@
III Single crystal neutron diffraction
III.1 Experiment and refinement details
@@
To determine the magnetic structure, two models were tested using the dataset obtained at 90 K. The first model adopts the magnetic space group , where the magnetic moments align along the crystallographic axis. The second model corresponds to the magnetic space group , in which the moments are oriented along the reciprocal axis. Although the model yielded a slightly better refinement result, the model is more consistent with our previous neutron diffraction results under strain-induced detwinning of magnetic domains [liu_StraintunableAnomalousHall_2025].
Table S3: Li-doped -MnTe at 340 K. Models placing Li at Wyckoff site , site , or uncorrelated interstitial sites were tested.
@@
We then tested alternative models incorporating Li atoms at crystallographic sites. Both the 2 and 2 positions (the latter coinciding with Te sites) yielded stable refinements. However, when Li atoms were placed at the 2 site, the refined concentration was significantly lower than the nominal doping level of 5%, as shown in Table S9. A similar analysis using the 90 K dataset produced negative Li occupancies at the 2 site when the atomic displacement parameters (ADPs) were not constrained, which is physically implausible.
@@
In accordance with previous powder neutron diffraction studies [moseley_giant_2022], we applied the orthorhombic magnetic space group to refine the magnetic structure of Li-doped -MnTe. This model also yielded a vanishingly small in-plane magnetic moment at the given doping level. In addition, we tested a hexagonal magnetic space group with higher symmetry, in which the magnetic structure is strictly antiferromagnetic and the magnetic moments are constrained to lie along the crystallographic axis. The refinement results for both models are summarized in Table S4.
III.2 Fractional atomic coordinates and displacement parameter (Å2) determined from neutron diffraction
@@
IV Survey of magnon linewidth across the Brillouin zone
Figure S2: (Top) Full-width at half-maximum (FWHM) values obtained from Gaussian fits of the spin wave spectra in the nodal planes of the Brillouin zone for undoped -MnTe and Li-doped -MnTe. Dashed lines indicate the energy-dependent instrumental resolution from Fig. S1. (Bottom) FWHM values obtained by subtracting the FWHM of undoped -MnTe from that of Li-doped -MnTe. Vertical error bars represent 1 standard deviations and horizontal error bars indicate the integration range for each line cut.
V INS data for Li-doped -MnTe taken using the HR mode
@@
VI Additional transport data
@@
VII Local moment sum rule
@@
is the dynamic contribution where is the -averaged local dynamic susceptibility, , and is the Boltzmann constant [RevModPhys.87.855].
@@
where is assumed to be in absolute units of b sr-1 f.u.-1 meV-1, 1 b = 10-24 cm2 is the unit “barn” for the neutron scattering cross section, is the magnetic form factor, and is the Debye-Waller factor [10.1063/1.4818323].
@@
VIII Momentum- and energy-integration ranges for the data
In the colormaps Fig. 2(a)(b), S2 and S3, the integration and plotting directions are mutually perpendicular. The integration ranges were chosen to ensure that the integrated volume in reciprocal space (in units of Å-1) is consistent across different directions throughout each figure. Integration ranges for other figures are provided in Table S12. For datasets where altermagnetic splitting is expected, such small integration ranges may be necessary to resolve the splitting clearly.
Table S11: Momentum-integration ranges for Fig. 2 (a)-(b), S2 and S3
Plot direction
Integration
direction 1
Integration
range 1 (r.l.u.)
Integration
direction 2
Integration
range 2 (r.l.u.)
@@ VIII Momentum- and energy-integration ranges for the data
[, -K, 0]
0.05
[0, 0, ]
0.093

### v2

@@ I.3 Inelastic neutron scattering
INS measurements on Li-doped -MnTe were carried out using the TOF spectrometer MAPS at the ISIS Neutron and Muon Source, Rutherford Appleton Laboratory [23]. The sample, with a mass of approximately 2.2 g, was obtained from the same large single crystal used for the TOPAZ experiment and was mounted in the [, 0, ] scattering plane. Data were collected at K with an incident energy of meV under continuous rotation about the vertical axis of the sample. During the measurements, the accelerator operated with a synchrotron current of 152 A on target station 1. Measurements were carried out in two modes: high-resolution (HR) mode with a Fermi chopper frequency of Hz and a sample rotation range of 60 degrees, and high-flux (HF) mode with Hz and a rotation range of 120 degrees. Each mode was allocated approximately 60 hours of beamtime, evenly distributed over the full angle range. The data were reduced into .nxspe files using a step of 0.5 degrees. Data from the HF mode are presented in Figs. 2(a)(c)–(e), 3(b)(d)(f)–(i), 4, and S2, while data from the HR mode are shown in Figs. 2(c)–(e), S3, and S4.
Data reduction was performed using MANTID [3], Horace [22], and SHIVER [1].
@@
I.4 Absolute unit normalization
INS data were normalized to absolute units during the data reduction process by measuring a standard vanadium sample of known mass [72]. The vanadium measurements were performed in the same sample environment and with the same chopper and instrument settings as the sample measurements.
I.5 Linear spin wave theory (LSWT) Calculations
The spin Hamiltonian was determined from the INS spectra using linear spin wave theory (LSWT) implemented in the SpinW package [70]. For both compounds, a Heisenberg spin Hamiltonian with a single-ion anisotropy term was employed, expressed as
where indexes the Mn2+ ions, is the spin operator at site , and denotes its component. represents the exchange interaction between spins at sites and , with the summation taken over spin pairs. In this coordinate system, the -axis corresponds to the crystallographic -axis. The anisotropy term is positive for easy-plane anisotropy, as in the case of pure -MnTe, and negative for Ising-type anisotropy, as observed in Li-doped -MnTe.
To extract the exchange coupling parameters using the SpinW program,
we first constructed the lattice and magnetic structures and enabled the
relevant exchange interaction and single-ion anisotropy terms.
The lattice and magnetic structures were obtained from single-crystal neutron
diffraction results and CIF files available in the ICSD database.
The extracted spin-wave dispersion relation was then used
as input for the SpinW calculations. To determine , we performed Gaussian fits to the inelastic neutron
scattering spectra at each wave vector along high-symmetry directions.
From these fits, we extracted the peak center energy and the half-width at
half-maximum (HWHM).
For the models “Pure -MnTe center” and “Li-doped -MnTe center”
listed in Table S1, the fitted peak energies at each
along the high-symmetry directions were directly used in the SpinW
calculations. Unless otherwise specified, all spin wave calculations presented in the main text for Li-doped MnTe are based on the exchange coupling parameters obtained from “Li-doped -MnTe center” model.
For the models “Li-doped -MnTe upper” and “Li-doped
-MnTe lower,” we introduced a modification to account for
spin wave broadening. At the Brillouin zone (BZ) boundaries (i.e.,
along the M–L and L–A paths), we used and
, respectively, to represent the upper and lower
bounds of the broadened dispersion. Along all other high-symmetry directions,
we used the unmodified fitted peak energy , since the apparent broadening there is mainly
influenced by the degeneracy of optical and acoustic modes.
@@ I.5 Linear spin wave theory (LSWT) Calculations
I.6 Electrical transport measurements
@@
All ARPES data were taken at beamline 7.0.2 of the Advanced Light Source at Lawrence Berkeley National Lab using the deflector mode of a Scienta R4000 analyzer with an energy resolution of less than 20 meV. The samples were cleaved in-situ with a base pressure better than Torr. All measurements were performed at a temperature of 15 K using 120 eV photons with linear horizontal polarization.
II Single crystal neutron diffraction
II.1 Experiment and refinement details
@@
For both -MnTe and Li-doped
-MnTe, we find no evidence of any other impurity phases, clearly different from powder samples of Li-doped -MnTe [73].
@@
To determine the magnetic structure, two models were tested using the dataset obtained at 90 K. The first model adopts the magnetic space group , where the magnetic moments align along the crystallographic axis. The second model corresponds to the magnetic space group , in which the moments are oriented along the reciprocal axis. Although the model yielded a slightly better refinement result, the model is more consistent with our previous neutron diffraction results under strain-induced detwinning of magnetic domains [44].
Table S3: Li-doped -MnTe at 340 K. Models placing Li at Wyckoff site , site , site , or uncorrelated interstitial sites were tested.
@@
We then tested alternative structural models incorporating Li atoms at
different crystallographic sites. Stable refinements were obtained when
Li was placed at the , , and positions (the latter coinciding
with the Te site). However, when Li atoms were assigned to the site,
the refined concentration was significantly lower than the nominal doping
level of 5%, as shown in Table S10. A similar refinement using the 90 K dataset yielded negative Li occupancies
at the site when the atomic displacement parameters (ADPs) were not
constrained, which is physically implausible. Placing Li at the site
resulted in a refined concentration of 1.7%, still much smaller than
the nominal doping level. Moreover, refinement of the 340 K dataset in this
model produced unphysical negative anisotropic ADPs. Therefore, we conclude that neither the nor the site provides a
physically reasonable model to account for the observed neutron diffraction
intensities.
@@
In accordance with previous powder neutron diffraction studies [48], we applied the orthorhombic magnetic space group to refine the magnetic structure of Li-doped -MnTe. This model also yielded a vanishingly small in-plane magnetic moment at the given doping level. In addition, we tested a hexagonal magnetic space group with higher symmetry, in which the magnetic structure is strictly antiferromagnetic and the magnetic moments are constrained to lie along the crystallographic axis. The refinement results for both models are summarized in Table S4.
II.2 Fractional atomic coordinates and displacement parameter (Å2) determined from neutron diffraction
@@
Table S10: Li-doped -MnTe, 340 K with the assumption at Li atoms sit at the 4 site
Atom
Label
Occ.
Site
Sym.
Mn
Mn1
0
0
0
1.000
0.018
2
Te
Te1
1/3
2/3
1/4
0.959(3)
0.010
2
Li
Li1
1/3
2/3
0.898(14)
0.009(6)
0.02
4
@@
III INS data for Li-doped -MnTe taken using the HR mode
@@
IV Energy-dependence of magnon lifetime in Li-doped -MnTe
Figure S4: (a) Energy-dependence of the full width at half-maximum (FWHM) for Li-doped -MnTe in high-flux (HF) mode (filled symbols) and high-resolution (HR) mode (open symbols). Dashed line serves as a guide to the eye marking the onset of increased linewidth above meV. (b) Differences in FWHM (FWHM) between Li-doped -MnTe and -MnTe along various high-symmetry directions in the nodal planes of the BZ. The top panel utilizes HF data for both Li-doped -MnTe and -MnTe, where the resolution is comparable between Li-doped -MnTe and -MnTe measurements. The bottom panel uses HR data for Li-doped -MnTe and HF data for -MnTe, meaning the resolutions are different for each.
FWHM values were subtracted at identical for both samples, and the corresponding energy of the Li-doped -MnTe dataset () is used as the -axis value for each point.
V Additional transport data
@@
VI Local moment sum rule
@@
is the dynamic contribution where is the -averaged local dynamic susceptibility, , and is the Boltzmann constant [16].
@@
where is assumed to be in absolute units of b sr-1 f.u.-1 meV-1, 1 b = 10-24 cm2 is the unit “barn” for the neutron scattering cross section, is the magnetic form factor, and is the Debye-Waller factor [72].
@@
VII Momentum- and energy-integration ranges for the data
In Figures 2, S2, and S3, the momentum integration directions and the plotting direction (if applicable) are mutually perpendicular. The integration ranges were chosen to ensure that the integrated volume in reciprocal space (in units of Å-1) is consistent across different directions throughout each figure. Integration ranges for other figures are provided in Table S14. For datasets where altermagnetic splitting is expected, such small integration ranges may be necessary to resolve the splitting clearly.
Table S12: Momentum-integration ranges for Fig. 2 (a)-(b) and S2
Path
Momentum 1 (r.l.u)
Momentum 2 (r.l.u)
@@ VII Momentum- and energy-integration ranges for the data
[, , 0], 0.05
[0, 0, ], 0.093
Table S13: Momentum-integration ranges for Fig. 2(c)-(e)
Path
Momentum 1 (r.l.u)
Momentum 2 (r.l.u)
Momentum 3 (r.l.u)
K-
[, , 0], 0.04
[0, , 0], 0.03
[0, 0, ], 0.06
-M
[, 0, 0], 0.03
[, , 0], 0.03
[0, 0, ], 0.06
M-L
[0, 0, ], 0.06
[, , 0], 0.03
[0, , 0], 0.03
L-A
[, 0, 0], 0.03
[, , 0], 0.03
[0, 0, ], 0.06
-A
[0, 0, ], 0.06
[, , 0], 0.03
[0, , 0], 0.03
-H
[, , 0], 0.04
[0, , 0], 0.03
[, , ], 0.009